Find the measure of each exterior angle of a regular polygon of 
15 sides

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of angles a regular polygon having side 15 = (15-2)×180° = 13×180° = 2340°
Each interior angle = 2340/15 = 156°
Each exterior angle = 180° - 156° = 24°
Or,
Each exterior angle = sum of exterior angles/Number of angles = 360/15 = 24°

Thanks